Find two consecutive positive even integers such that the square of the second, decreased by twice the first is 124 find the numbers.
Show steps and use let statements

Answer by ikleyn(52790)

.

Seeking for the numbers n, n+2, you have this equation


    (n+2)^2 - 2n = 124


    n^2 + 2n + 4 = 124


    n^2 + 2n = 120

   
    n*(n+2) = 120


    n = 10               ANSWER

Solved.